# The unreasonable likelihood of being Origin of life, terraforming, and AI

### v1

@@ Abstract
The genomic information is the easiest to estimate—and, perhaps surprisingly, contributes the least. Consider a minimal self-replicating organism such as Mycoplasma genitalium, which contains roughly 500 genes [37], each about 1,000 base pairs long. With two bits per base pair (for A, T, C, G), this yields $I_{g}\sim 500\times 1000\times 2=10^{6}$ bits. This estimate is supported by the synthetic minimal cell JCVI-syn3.0, developed from Mycoplasma mycoides, which contains only 473 genes [51]. Remarkably, 149 of these genes still have unknown function—underscoring how much we have yet to understand, even in minimal life.
@@
To estimate $I_{s}$, we consider protein structure and folding. A protein’s biological function depends on its shape. While AlphaFold2—an AI-based protein-folding algorithm—occupies 556 GB, this includes thousands of structures, redundancies, and metadata [35]. A more realistic lower bound for a minimal cell might be in the range of $10^{6}$$10^{8}$ bits, depending on the number of unique folds, protein domains, and the information needed to describe each.
@@
To estimate $I_{d}$, the dynamic component, we turn to whole-cell models. The first comprehensive whole-cell simulation of Mycoplasma genitalium integrated 28 submodels covering DNA replication, transcription, translation, metabolism, and more [37]. The compressed code size was about 140 MB, or roughly $10^{9}$ bits. The model used over 1,900 parameters curated from 900+ publications and could predict phenotypes from genotypes. While this is almost certainly an overestimate—due to redundant code and parameters—it’s informative. Notably, the model excludes spatial organization, mechanical feedback, and memory effects. Other whole-cell simulations of E. coli [38] and synthetic cell JCVI-syn3A [52] use somewhat smaller codebases but follow similar principles.
@@
corresponding to an accumulation rate of only about 2 bits per year.
Compared to $R(D)\sim 100$ bits/s (modulo the efficiency factor $\eta$), this is minuscule—but reflects the vast timescale over which life could have emerged. To illustrate: a DNA molecule grows by 2 bits per base pair. At a rate of 100 bits/s, this corresponds to 50 base pairs per second, or more than the considerable $500$ m length of a DNA molecule in a year. The remarkably low threshold for information accumulation at $6.34\times 10^{-8}$ bits/s suggests that abiotic emergence of life should be feasible, even with extremely low efficiency ($\eta\sim 10^{-8}$) or short molecular lifetimes ($D\sim 1$ s).
Drift versus random diffusive process for self-assembly
The minimum information accumulation rate required to assemble a protocell of complexity $I_{\text{protocell}}=10^{9}$ bits within $T_{\text{available}}=500$ million years is $v_{\text{min}}=R_{\text{min}}=6.34\times 10^{-8}$ bits/s, highlighting that even extremely inefficient or sporadic processes could, in principle, accumulate the required information over geological timescales. However, this estimate assumes a unidirectional, steadily progressive (ballistic or forward-biased) process. What if the information-gathering process behaves more like a random walk—with fluctuations and reversals—rather than directed drift?
Let us consider a stochastic accumulation model: a biased random walk with velocity $v$ and persistence time $\tau$, analogous to “run-and-tumble” dynamics from bacterial chemotaxis [53]. The effective diffusion constant is given by $D_{\text{eff}}=v^{2}\tau/d$, where $d$ is the dimensionality of the search space. If $v=v_{\text{min}}$, then the mean-square displacement (MSD) after time $T_{\text{available}}$ is
$\langle I^{2}_{\text{protocell}}\rangle=(v_{\text{min}}T_{\text{available}})^{2}=2dD_{\text{eff}}T_{\text{available}}=2v_{\text{min}}^{2}\tau T_{\text{available}},$
independent of $d$. Solving for the crossover time $\tau_{c}$ at which this diffusive process accumulates the same information as directed drift yields $\tau_{c}=T_{\text{available}}/2$. Thus, if $\tau<\tau_{c}=250$ Myr, the randomization is too strong, and the process fails to reach the target within $T_{\text{available}}$ (Fig. 3). High persistence—i.e., strong directional memory—is essential.
To illustrate the timescale sensitivity: With $\tau=1$ s, the wait time for assembling a cell becomes $10^{24}$ yr — roughly a hundred trillion universes stacked end to end. Even if $\tau=1$ yr, the required time is still $T^{\prime}\approx 10^{17}$ yr, about ten million times the universe’s current age. In other words, without immense persistence, life’s emergence becomes cosmologically implausible, potentially pointing to alternative mechanisms.
In summary, the plausibility of information accumulation via biased random walks hinges critically on both the persistence time $\tau$ and the effective information accumulation rate $v$. Highly biased or driven processes dramatically outperform fully diffusive ones in crossing the informational threshold required for life.
Figure 3: How driven does a random walk need to be in information space? Time to reach protocell complexity $I_{\text{protocell}}$ as a function of the persistence time $\tau$ for a biased random walk using minimal information accumulation speed $R_{\text{min}}$, given by $T^{\prime}=\langle I_{\text{protocell}}\rangle^{2}/(2v_{\text{min}}^{2}\,\tau)$ (blue line). The available time of $500$ Myr (red dashed line) is reached for a persistence time of $250$ Myr. Also shown is the age of the universe (yellow line).
Material uptake and energy consumption are not limiting
[…]
Using the minimum required information rate $v_{\text{min}}\approx 6.34\times 10^{-8}$ bits/s and again assuming 10 bits per useful molecule, the corresponding capture rate is $v_{\text{min}}/10\approx 6.34\times 10^{-9}$ molecules/s. This implies that, on average, fewer than one molecule needs to be captured every 5 years. Compared to the Berg and Purcell limit of $10^{9}$ molecules/s, this rate is 13 orders of magnitude lower. Hence, even extremely sparse and selective molecular capture would suffice to meet the informational demands of life’s emergence, given enough time.
If a protocell requires $I_{\text{protocell}}\sim 10^{9}$ bits, and each molecule contributes approximately 10 bits, then the total number of molecules needed is $N_{\text{protocell}}=I_{\text{protocell}}/10=10^{8}$ molecules, consistent with estimates for E. coli [55]. Assuming diffusion-limited uptake, the time required to accumulate $N_{\text{protocell}}$ molecules is $\Delta t=N_{\text{protocell}}/J=10^{8}/(2.27\times 10^{9})\approx 0.044\ \text{s}\approx 1.4\times 10^{-9}$ yr. In principle, a protocell could acquire all necessary components in under 50 milliseconds—if every incoming molecule were retained and useful. In reality, however, most molecules would be non-functional, degraded, or unincorporated, making molecular selection and retention—not flux—the true bottlenecks in prebiotic information accumulation.
To estimate the energy required to assemble a system with $I_{\text{protocell}}\sim 10^{9}$ bits, we compare three frameworks: the Landauer limit, biological ATP-driven synthesis, and technological processes such as PCR. First, the Landauer principle states that erasing one bit of information incurs a minimum energy cost of $E_{\text{min}}=k_{B}T\ln 2\approx 2.8\times 10^{-21}$ J/bit at $T\approx 300$ K. Thus, the theoretical minimum energy for assembling the protocell is $E_{\text{Landauer}}\approx 10^{9}\times 2.8\times 10^{-21}=2.8\times 10^{-12}\ \text{J}$. Second, biological energy costs based on ATP hydrolysis provide about $10^{-19}$ J per molecule. If synthesizing one informational molecule (containing $\sim$10 bits) requires one ATP (a conservative assumption), then $E_{\text{ATP}}=10^{8}\times 10^{-19}=10^{-11}\ \text{J}$. Third, technological synthesis via polymerase chain reaction (PCR) consumes about 1–10 J to amplify microgram-scale DNA. A 109-bit construct corresponds to $\sim$50 $\mu$g of DNA, implying $E_{\text{PCR}}\sim 1\ \text{J}$, or roughly $10^{-9}$ J/bit. Even 1 J is a small amount of energy—equivalent to lifting a 100 g apple by 1 meter or powering a smartphone for 1 second.
Hence, minimal life appears to operate within one to two orders of magnitude of the thermodynamic limit, while modern technology (e.g. PCR) is up to $10^{9}$ times less efficient. This contrast highlights the remarkable energetic efficiency of biological systems and provides a thermodynamic baseline for evaluating the feasibility of prebiotic assembly.
@@
Above, we estimated the information content of a minimal protocell using Kolmogorov complexity, based on detailed quasi-realistic descriptions of cellular processes encoded in modern computational models. This supports a picture of life emerging through a gradual, steady increase in the organization and complexity of matter—from nonliving to living. This is like putting together a gigantic jigsaw puzzle, and appears like the culmination of a reductionistic belief system. And the required persistence times are extreme, even mind boggling.

### v2

@@ Abstract
The genomic information is the easiest to estimate—and, perhaps surprisingly, contributes the least. Consider a minimal self-replicating organism such as Mycoplasma genitalium, which contains roughly 500 genes [Karr2012], each about 1,000 base pairs long. With two bits per base pair (for A, T, C, G), this yields $I_{g}\sim 500\times 1000\times 2=10^{6}$ bits. This estimate is supported by the synthetic minimal cell JCVI-syn3.0, developed from Mycoplasma mycoides, which contains only 473 genes [Hutchison2016]. Remarkably, 149 of these genes still have unknown function—underscoring how much we have yet to understand, even in minimal life.
@@
To estimate $I_{s}$, we consider protein structure and folding. A protein’s biological function depends on its shape. While AlphaFold2—an AI-based protein-folding algorithm—occupies 556 GB, this includes thousands of structures, redundancies, and metadata [Jumper2021]. A more realistic lower bound for a minimal cell might be in the range of $10^{6}$$10^{8}$ bits, depending on the number of unique folds, protein domains, and the information needed to describe each.
@@
To estimate $I_{d}$, the dynamic component, we turn to whole-cell models. The first comprehensive whole-cell simulation of Mycoplasma genitalium integrated 28 submodels covering DNA replication, transcription, translation, metabolism, and more [Karr2012]. The compressed code size was about 140 MB, or roughly $10^{9}$ bits. The model used over 1,900 parameters curated from 900+ publications and could predict phenotypes from genotypes. While this is almost certainly an overestimate—due to redundant code and parameters—it’s informative. Notably, the model excludes spatial organization, mechanical feedback, and memory effects. Other whole-cell simulations of E. coli [AhnHorst2022] and synthetic cell JCVI-syn3A [Domenzain2022] use somewhat smaller codebases but follow similar principles.
We now turn to $I_{a}$, the assembly component of the information required to make a protocell, and ask how unlikely it is for such an object to arise through random chemical assembly alone. Instead of invoking more elaborate assembly theories, we use a simple branching-process picture. We consider an object of effective assembly length $L$, with a branching factor $b$ describing the number of random chemical choices available at each step. The total number of possible assembly paths then scales as $N_{\text{paths}}\sim b^{L}$, while the number of paths that successfully produce the target object is small, $N_{\text{object}}$, plausibly scaling only polynomially with $L$, $N_{\text{object}}\sim\text{poly}(L)$, or even $\mathcal{O}(1)$ as a first approximation.
The probability that a random assembly trajectory of length $L$ reaches the target structure is therefore
$P_{\text{path}}(X)\sim\frac{L^{K}}{b^{L}}\approx b^{-L},$ (4)
up to polynomial prefactors. This expression formalizes the intuition that for complex objects of large size $L$, the set of successful assembly histories forms a vanishingly thin corridor within an exponentially large tree of possibilities. Taking the negative logarithm of this probability yields a natural estimate of the assembly information cost associated with realizing the target structure. This argument can be extended to $M$ modules assembled hierarchically. [^2] For a complex structure composed of $M$ modules, one may write $P_{\text{path}}(X)\sim\sum_{i}^{M}L_{i}^{K}/b_{i}^{L_{i}}+(\text{joining terms})$ for modules of size $L_{i}$ and branching factor $b_{i}$, with $i=1,\ldots,M$.
Plugging in some numbers for an order of magnitude estimate, we may say there are $M\sim 10^{3}$ modules, each typically made from $L\sim 10^{3}$ key effective steps, and an effective branching factor of $b_{\text{eff}}\sim 10-10^{3}$, describing possible reaction or attachments in the prebiotic chemical environment. Neglecting joining issues from the different modules for simplicity, this would lead to $I_{a}\sim ML\log_{2}(b_{\text{eff}})\sim 10^{6}\times(3-10)$ bits $\sim 10^{7}$ bits, but, due to the immense difficulty in estimating this, could equally lead to larger numbers like $10^{9}$ bits similar to $I_{s}$ or even larger, forming the bottleneck of protocell formation. Hence, a cell is not only informationally complex in its final configuration; it also sits at the end of an astonishingly long and narrow assembly corridor in process space.
@@
$R_{\text{min}}=\frac{I_{\text{protocell}}}{T_{\text{available}}}\approx 6.34\times 10^{-8}\ \text{bit/s},$ (5)
corresponding to an accumulation rate of only about 2 bits per year. Compared to $R(D)\sim 100$ bit/s (modulo the efficiency factor $\eta$), this is minuscule—but reflects the vast timescale over which life could have emerged. To illustrate: a DNA molecule grows by 2 bits per base pair. At a rate of 100 bit/s, this corresponds to 50 base pairs per second, or more than the considerable $500$ m length of a DNA molecule in a year. The remarkably low threshold for information accumulation at $6.34\times 10^{-8}$ bit/s suggests that abiotic emergence of life should be feasible, even with extremely low efficiency ($\eta\sim 10^{-8}$) or short molecular lifetimes ($D\sim 1$ s).
Drift versus random diffusive process for self-assembly
The minimum information accumulation rate required to assemble a protocell of complexity $I_{\text{protocell}}=10^{9}$ bits within $T_{\text{available}}=500$ million years is $v_{\text{min}}=R_{\text{min}}=6.34\times 10^{-8}$ bit/s, highlighting that even extremely inefficient or sporadic processes could, in principle, accumulate the required information over geological timescales. However, this estimate assumes a unidirectional, steadily progressive (ballistic or forward-biased) process. What if the information-gathering process behaves more like a random walk—with fluctuations and reversals—rather than directed drift?
Let us consider a stochastic accumulation model: a random walk with velocity $v$ and persistence time $\tau$, analogous to “run-and-tumble” dynamics from bacterial chemotaxis [Berg1993]. The effective diffusion constant is given by $D_{\text{eff}}=v^{2}\tau/d$, where $d$ is the dimensionality of the search space. If $v=v_{\text{min}}$, then the mean-square displacement (MSD) after time $T_{\text{available}}$ is
$\langle I^{2}_{\text{protocell}}\rangle=(v_{\text{min}}T_{\text{available}})^{2}=2dD_{\text{eff}}T_{\text{available}}=2v_{\text{min}}^{2}\tau T_{\text{available}},$ (6)
independent of $d$. Solving for the crossover time $\tau_{c}$ at which this diffusive process accumulates the same information as directed drift yields $\tau_{c}=T_{\text{available}}/2$. Thus, if $\tau<\tau_{c}=250$ Myr, the randomization is too strong, and the process fails to reach the target within $T_{\text{available}}$ (see Fig. 3A). High persistence—i.e., strong directional memory—is essential. To illustrate the timescale sensitivity: With $\tau=1$ s, the wait time for assembling a cell becomes $10^{24}$ yr — roughly a hundred trillion universes stacked end to end. Even if $\tau=1$ yr, the required time is still $T^{\prime}\approx 10^{17}$ yr, about ten million times the universe’s current age. In other words, without immense persistence, life’s emergence becomes cosmologically implausible, potentially pointing to alternative mechanisms.
However, this result may be too restrictive as we solely allowed minimum information accumulation speed $v_{\text{min}}$ which is determined to make the drift scenario just work in the available time $T_{\text{available}}$. A persistence time of $\tau_{c}=T_{\text{available}}/2$ simply reflects that, after doing one initial step towards a protocell, the next step, another gigantic ”fluctuation” of random organization over $T_{\text{available}}/2$, either completes the protocell, or undoes any progress made, leading to a completed protocell with 50% chance. Let’s allow $v>v_{\text{min}}$ and hence $\tau_{c}^{\prime}<\tau_{c}$ via Eq. 6 with $v_{\text{min}}$ replaced by $v$ on the right-hand side and solving for $\tau=\tau_{c}^{\prime}$. Now, $\tau_{c}^{\prime}=(v_{\text{min}}/v)^{2}T_{\text{available}}/2$. The result is shown in Fig. 3B. Making a protocell in $T_{\text{available}}$ with a persistence time of e.g. ”only” $\tau_{c}^{\prime}=1,000$ yr now requires $v=500\times v_{\text{min}}\approx 3.2\cdot 10^{-5}$ bit/s (or equivalently of about $500$ bp/yr), and for $\tau_{c}^{\prime}=1$ s (still large for molecular reaction time scales) now requires $v\approx 6$ bit/s (or $2.8$ bp/s). Hence, while some of those persistent times and speeds seem reasonable to make a protocell by random walk, this has to be done consistently for $500$ Gyr, even though the climate is fluctuating, especially on the early planet Earth.
Up to now we laid out the main contribution of this paper: the emergence of a protocell is possible, but only if information accumulation via random walks has a large persistence time (or a high effective information accumulation rate), over extended (enormous) periods of time. However, the devil is in the details, specifically factor $\eta$ - the actual physio-chemical mechanism and the conceptual big unknown. This factor could be close to zero, rendering the rate–distortion bound unattainable in practice. In the following we examine several factors influencing $\eta$, including energetic requirements (which appear comparatively permissive), assembly kinetics (which likely constitute the dominant bottleneck), and nonequilibrium phase transitions (which may offer a potential route to overcoming these limitations).
Figure 3: How persistent does a random walk need to be in information space? (A) Time to reach protocell complexity $I_{\text{protocell}}$ as a function of the persistence time $\tau$ for a random walk using minimal information accumulation speed $R_{\text{min}}$, given by $T^{\prime}=\langle I_{\text{protocell}}^{2}\rangle/(2v_{\text{min}}^{2}\,\tau)$ (blue line). The available time of $500$ Myr (red dashed line) is reached for a persistence time of $250$ Myr. Also shown is the age of the universe (yellow line). (B) Persistence time for allowing larger information accumulation speeds than $v_{\text{min}}$ for protocell formation in $T_{\text{available}}$. The larger the speed of information accumulation, the shorter the required persistence time of the random walk. For details see the github repository.
Material uptake and energy consumption are not limiting
To assess whether the prebiotic environment could support the molecular uptake rate required by our information-theoretic model, we compare it to Berg and Purcell’s classical estimate for the molecular flux to a spherical cell [BergPurcell1977]. For a cell of radius $a=1\,\mu\mathrm{m}$, diffusion coefficient $D=300\,\mu\mathrm{m}^{2}/\mathrm{s}$, and concentration $c=1\,\mathrm{mM}$, the diffusive flux is given by $J=4\pi Dac\approx 2.27\times 10^{9}\,\text{molecules/s}$.
In contrast, our rate-distortion estimate requires an information accumulation rate of $R(D)\sim 100\,\text{bit/s}$, with each molecule contributing approximately 10 bits. This implies a necessary uptake of only $R(D)/10=10$ molecules/s. The actual molecular flux thus exceeds the required rate by a factor of approximately $10^{8}$. Therefore, the prebiotic chemical environment was not transport-limited—even at nanomolar concentrations. Rather, the challenge lay in filtering, stabilizing, and retaining the few informationally useful molecules from the overwhelming molecular background.
Using the minimum required information rate $v_{\text{min}}\approx 6.34\times 10^{-8}$ bit/s and again assuming 10 bits per useful molecule, the corresponding capture rate is $v_{\text{min}}/10\approx 6.34\times 10^{-9}$ molecules/s. This implies that, on average, fewer than one molecule needs to be captured every 5 years. Compared to the Berg and Purcell limit of $10^{9}$ molecules/s, this rate is 13 orders of magnitude lower. Hence, even extremely sparse and selective molecular capture would suffice to meet the informational demands of life’s emergence, given enough time.
If a protocell requires $I_{\text{protocell}}\sim 10^{9}$ bits, and each molecule contributes approximately 10 bits, then the total number of molecules needed is $N_{\text{protocell}}=I_{\text{protocell}}/10=10^{8}$ molecules, consistent with estimates for E. coli [Milo2015CellBiologyByNumbers]. Assuming diffusion-limited uptake, the time required to accumulate $N_{\text{protocell}}$ molecules is $\Delta t=N_{\text{protocell}}/J=10^{8}/(2.27\times 10^{9})\approx 0.044\ \text{s}\approx 1.4\times 10^{-9}$ yr. In principle, a protocell could acquire all necessary components in under 50 milliseconds—if every incoming molecule were retained and useful. In reality, however, most molecules would be non-functional, degraded, or unincorporated, making molecular selection and retention—not flux—the true bottlenecks in prebiotic information accumulation.
@@
As a formal aside, and solely to obtain a mechanistic lower bound on energetic cost, it is useful to frame the problem in terms of unbalanced optimal transport [UOT2016]. Let the environment contain molecular building-blocks of types $i=1,\dots,B$ with abundances $\mu_{i}$, and let a protocell require molecular species $j=1,\dots,N$ in quantities $\nu_{j}$. A transport–transformation plan $\pi_{ij}$ specifies how many molecules of type $i$ are converted into molecules of type $j$. The unbalanced case is appropriate here because the environment generally does not contain enough of the specific molecules needed; creation of new molecules from simpler precursors is therefore required. A simple unbalanced optimal transport cost functional is
$\mathrm{cost}(\pi)=\sum_{i,j}c_{ij}\,\pi_{ij}\;+\;\kappa_{\mathrm{tgt}}\sum_{j}\left(\nu_{j}-\sum_{i}\pi_{ij}\right)_{+},$ (7)
where $c_{ij}$ denotes the free-energy cost of transforming a molecule of type $i$ into one of type $j$, and $\kappa_{\mathrm{tgt}}$ is the free energy required to synthesize a molecule of type $j$ “from scratch” if the environment does not supply enough precursors. The $(\cdot)_{+}$ denotes the positive part, ensuring that shortages must be paid for by de novo synthesis. This functional has a direct physical interpretation: it computes the minimal reversible Gibbs free energy required to convert an environmental composition vector $\mu$ into the target protocell composition $\nu$, assuming perfectly efficient chemistry, perfect selectivity, and no kinetic or spatial constraints.
For present purposes it is useful to make the most conservative assumption possible, namely that the environment does not conveniently provide large amounts of completed lipids, nucleotides, or peptides. In this case virtually all material must be created by de novo synthesis, and the optimisation of (7) collapses to a sum over required protocell components,
$E_{\min}\simeq N_{L}\,\Delta G_{L}+N_{N}\,\Delta G_{N}+N_{P}\,\Delta G_{P},$ (8)
where $_{L}, NN$$_{N}$, and $N_{P}$ are the numbers of membrane lipids, nucleotide monomers, and amino-acid residues required, respectively. The quantities $\Delta G_{L}$, $\Delta G_{N}$, and $\Delta G_{P}$ represent lower-bound reversible free energies for synthesising each corresponding molecule from simple bulk precursors (CO2, H2, N2, etc.). These free-energy values implicitly include the minimal chemical work, the minimal transport work (such as insertion of lipids into a membrane), and the minimal assembly work (such as forming a phosphodiester or peptide bond), and are deliberately chosen to be small so that the resulting estimate is an extremely generous lower bound.
To obtain a numerical value, consider a toy protocell containing $N_{L}\sim 10^{8}$ lipids, $N_{N}\sim 10^{6}$ nucleotide monomers, and $N_{P}\sim 10^{6}$ amino-acid residues. Conservatively taking $\Delta G_{L}\sim 50\,\mathrm{kJ/mol}$, $\Delta G_{N}\sim 60\,\mathrm{kJ/mol}$, and $\Delta G_{P}\sim 40\,\mathrm{kJ/mol}$, and converting from molar to per-molecule scales, one finds
$E_{\min}\sim 10^{-11}\,\mathrm{J},$ (9)
which is already the most favourable reversible estimate obtainable from any physically meaningful choice of parameters. Expressed in thermal units at $T\approx 300\,\mathrm{K}$,
$\frac{E_{\min}}{k_{B}T}\sim 10^{9},$ (10)
and using the Landauer relation $k_{B}T\ln 2$ per bit, this corresponds to an effective thermodynamic complexity of approximately
$\frac{E_{\min}}{k_{B}T\ln 2}\sim 10^{9}\ \text{bits}.$ (11)
It is important to emphasise that this $E_{\min}$ represents only a lower bound on the minimal reversible work of assembly. It neglects all kinetic barriers, all dissipation, all side reactions, all issues of pathway search, all spatial transport limitations, all problems of chirality purification, and all requirements for templating or sequence order. Any realistic prebiotic scenario would require substantially more energy, but this most optimistic bound lies in the range suggested by independent arguments such as energy requirements ranging from $10^{-12}-10^{-9}$ J (or even $1$ J for modern inefficient PCR) and $I_{a}\sim 10^{7}-10^{9}$ bits. Hence, we are beginning to paint a consistent picture for evaluating the feasibility of prebiotic assembly.
Finding the assembly path is the tricky bit
Having argued that neither material flux nor energetic cost poses a fundamental obstacle, we now turn to what remains: the structure of assembly paths themselves. Earlier we used a simple branching process to estimate the complexity involved in assembling a protocell and then translated this into an energetic cost using the Landauer bound, yielding thermodynamic complexity estimates that are, in principle, manageable and consistent with earlier bounds. In reality, however, the assembly process is far more subtle. Molecules must form complex structures through hierarchical assembly in a manner that does not obstruct subsequent steps, and the entire process must function dynamically—as a machine in motion rather than a frozen sculpture.
To gain basic intuition, this assembly process is rather very different from assembling a 2D jigsaw puzzle. Even if a piece is missing in the middle, we can easily reach that space and fill it using the available third dimension; we can simply move it there, not affecting the remainder of the puzzle. The assembly path is simple and highly parallel: we can start in any corner, or anywhere in the middle, or do it in chunks and assemble the chunks later.
To formalize this difficult process, we turn to empirical assembly theory [Sharma2023Assembly] [^3] While a useful framework, ’theory’ is a strong word - it is more like an empirical metric of complexity., which is a framework that quantifies complexity by measuring the minimal number of assembly steps needed to build an object from its basic parts. It suggests that complex objects, especially those with biological origins, are unlikely to arise by random chance and require a specific, historical path to be built. The framework provides a measurable way to distinguish between objects made by random processes and those created through selection and evolution, and it has implications for detecting life on other planets.
In this framework, the assembly number $A$, is an empirical measure designed to quantify how strongly a molecule is supported by selection and replication versus passive abiotic chemistry [Sharma2023Assembly]. For a given molecular species $i$ with assembly index $a_{i}$ (the length of the shortest construction path in a restricted combinatorial space) and abundance $n_{i}$, they define
$A\;=\;\sum_{i=1}^{N}\exp(a_{i})\,\frac{n_{i}-1}{N_{T}},$ (12)
where $N_{T}=\sum_{i}n_{i}$ is the total number of molecules detectable above threshold. Although the expression resembles a Boltzmann weight, the factor $\exp(a_{i})$ is not thermodynamic; it is an empirical amplifier that gives exponentially greater weight to complex molecules (large $a_{i}$), while the factor $(n_{i}-1)/N_{T}$ encodes relative abundance. Thus $A$ is best viewed as a “complexity $\times$ abundance” score: large $A$ values indicate complex molecules that are also unexpectedly common relative to abiotic expectations, whereas small $A$ values correspond either too simple or too complex molecules whose abundance is negligible.
This structure predicts a characteristic bimodality when life is present. Abiotic chemistry alone is efficient at generating very low-assembly-index structures, but strongly suppresses high-assembly-index molecules. Under selection and replication, however, specific high-$a$ structures become amplified, creating a second peak in abundance. Between these two peaks, one generically expects a region of extremely low abundance—an “assembly desert”—because there is no smooth path of intermediate molecules that are both synthesized frequently enough and sufficiently stable. In other words, abiotic chemistry efficiently populates $a\approx 1$44; biology populates $a\gtrsim 20$; and the range in between is nearly empty.
In a simple toy model of only one key molecular species, Eq. 12 simplifies to $A\sim\exp(a)\;n$. This picture aligns with a simple branching-path model from earlier in which the probability of generating a molecule of assembly index $a$ by purely random, unguided exploration scales as $n_{\text{abiotic}}(a)\;\propto\;P_{\text{abiotic}}(a)\;\propto\;b^{-a}\,\mathrm{poly}(a)$, where $b$ is the average branching factor of the assembly graph, combined with a mild combinatorial prefactor of polynomial in $a$. The assembly index for possible structures of size $a$, we obtain
$A_{\text{abiotic}}(a)\;\sim\;\left(\frac{e}{b}\right)^{a}\,\mathrm{poly}(a).$ (13)
Since $b\gg e$ for any realistic molecular construction space, the factor $(e/b)^{a}$ decreases super-exponentially, so $A_{\text{abiotic}}(a)$ is sharply concentrated at very low $a$ and becomes effectively zero for modest $a\gtrsim 10$. This mathematically explains the expected emptiness (the “desert”) in the intermediate region.
In contrast, biology does not obey the branching process distribution. Once self-replication begins, abundances are driven by selection, not by combinatorics. For a molecular species replicated during metabolism or encoded by genes, $n_{\text{biotic}}(a)>\!\!>n_{\text{abiotic}}(a)$. Even species with very high assembly index may have many copies, produced repeatedly, stabilized, and protected from degradation. Plug that into the assembly index and we obtain:
$A_{\text{biotic}}(a)\;\sim\;e^{a}n_{\text{biotic}}(a).$ (14)
Even moderate amplification of complex molecules causes a huge increase in $A$ because $\exp(a)$ is already exponential. Combined with even a modest abundance $n_{\text{biotic}}(a)$, a second peak emerges. This is the “life signature.” Figure 4 shows results from a toy model, illustrating the two peaks for a single molecular species with assembly index $a$ and relative abundance $n$ (see caption for mathematical details).
The existence of such an empty bridging region is important for the origin-of-life problem. If there were a smooth, gradually decreasing abundance curve from low to high assembly index, then high-$a$ biomolecules could plausibly arise from incremental abiotic processes. Instead, both theory and observation suggest a discontinuity: a sparsely populated or empty interval separating simple molecules from the complex macromolecules that biology produces in abundance. This gap indicates that unguided prebiotic chemistry cannot efficiently traverse the exponentially expanding assembly space, and that only selection, templating, and replication can maintain a stable population of large-$a$ structures. The “assembly desert” therefore acts as a signature of life and simultaneously marks a fundamental barrier that abiotic chemistry must cross for life to emerge.
Unlike gaps in the fossil record, which can be explained by taphonomic biases and punctuated evolutionary equilibria [GouldEldredge1972Punctuated, EldredgeGould1997PEAge, BentonHarper2009Paleobiology, Prothero2007Fossils, Raup1976SpeciesDiversity] [^4] Speciation events require small, isolated populations, but fossilization requires large, stable, sediment-rich environments — the two rarely coincide., the absence of intermediate-complexity molecules in modern environments is a genuine chemical signal. Molecules do not fossilize, and if abiotic chemistry readily produced mid-complexity polymers, co-factors, nucleic-acid analogues, or protocell precursors, they should still occur today in hydrothermal vents, volcanic ponds, atmospheric photochemistry, or mineral interfaces [Benner2010DefiningLife, Sutherland2017EndBeginning, KitadaiMaruyama2018OriginsBlocks] [^5] Unless early Earth was very different from today’s extreme conditions, or complex molecules are constantly generated but we don’t look at the right place (e.g. environmental protective niches which can act as scaffolds of assembly), or we don’t recognize such complex molecule-breeding grounds. Most likely, however, the common sense answer is No - chemical complexity of intermediate level is not occurring anywhere at a high enough rate. . Yet modern Earth shows a stark bimodality: abundant simple molecules produced by geochemistry, and abundant complex molecules produced by life, with almost nothing in the middle [Lambert2008PrebioticPolymerization, DangerPlassonPascal2012Peptides, Cleaves2008PrebioticAminoAcids, PownerSutherland2008OnePotNucleotides, Bernhardt2012RNAWorldFragile]. This “assembly desert” cannot plausibly be explained by biological consumption, because cells do not import or metabolize arbitrary exotic polymers (e.g., PNA, TNA, random peptides) [LazcanoMiller1999Metabolism, KooninMartin2005OriginGenomes]. Rather, it reflects a genuine chemical bottleneck: the mid-complexity region is extremely difficult to populate without templating, compartmentalization, or sequence-controlled polymerization [Hazen2007emergence, Sasselov2020PlanetaryPhenomenon]. The absence of such molecules today is therefore evidence that the barrier they present is fundamental, and that early life must have required special mechanisms to cross it [CleavesBada2012AlternativeNucleicAcids].
Overall, assembly theory highlights a striking structural feature of chemical space: the deep gulf separating the simple, abundant molecules of abiotic chemistry from the complex, abundant molecules of biology. Its bimodal landscape makes clear that bridging this “assembly desert” requires mechanisms beyond trivial chemical exploration, sharpening—rather than resolving—the central puzzle of how life first crossed from one peak to the other. This reframes the origin of life not as a problem of energy or material supply, but as a problem of accessing and stabilizing rare assembly pathways.
Figure 4: Assembly theory illustrates the immense difficulty of assembling a protocell. (A) Assembly score from a toy model as a function of assembly index $a$ and relative abundance (copy number) $n$, for a single type of molecule (e.g. representing a bottleneck in assembly). There is a peak at low $a$ and high $n$ for abiotic chemistry, and another peak at high $a$ and intermediate $n$, representing a signature of life. To obtain such $a$ and $n$ values requires natural selection and growth. The white arrow indicates a kind of reaction coordinate connecting the two peaks. (B) Assembly score plotted along the reaction coordinate, illustrating the vast gap between the peaks. Mathematical details: plot shows $A(a,n)\propto e^{\alpha a}\;n\;(\rho_{\mathrm{abiotic}}+\rho_{\mathrm{life}})$, normalized to its maximum, based on an abiotic density $\rho_{\mathrm{abiotic}}(a,n)=b^{-a}\exp[-(n-n_{0}e^{-\gamma a})^{2}/(2\sigma_{n}^{2})]$ and a biotic Gaussian bump at high $a$ and moderate $n$, for suitable parameters $\alpha$, $b$, $n_{0}$, $\gamma$, $\sigma_{n}$ etc. For details see the github repository.
@@
Above, we estimated the information content of a minimal protocell using Kolmogorov complexity with some considerations of assembly, based on detailed quasi-realistic descriptions of cellular processes encoded in modern computational models. This supports a picture of life emerging through a gradual, steady increase in the organization and complexity of matter—from nonliving to living. This is like putting together a gigantic jigsaw puzzle, and appears like the culmination of a reductionistic belief system. And the required persistence times are extreme, even mind boggling.